### v1

# A lower bound on the right-handed neutrino mass from wash-in leptogenesis

Abstract
Leptogenesis is an attractive scenario for the generation of the baryon asymmetry of the Universe that relies on the dynamics of right-handed neutrinos (RHNs) in the seesaw extension of the Standard Model. In standard thermal leptogenesis, the RHN mass scale $M_{N}$ is subject to the Davidson–Ibarra bound, $M_{N}\gtrsim 10^{9}\,\textrm{GeV}$, which builds on the assumption that RHN decays are responsible for the violation of both charge-parity invariance ($CP$) and baryon-minus-lepton number ($B\!-\!L$). In this paper, we relax this assumption in the context of the more general framework of wash-in leptogenesis, in which $CP$ violation is encoded in the initial conditions and the only remaining task of the RHN decays is to violate $B\!-\!L$. Solving the relevant set of Boltzmann equations, we find that, in wash-in leptogenesis, the RHN mass scale can be as low as 7 TeV.
preprint: MITP-25-002preprint: MS-TP-24-01
Introduction — The baryon asymmetry of the Universe (BAU), commonly quantified in terms of the baryon-to-photon ratio $\eta_{B}^{\rm obs}=n_{B}^{0}/n_{\gamma}^{0}\simeq 6.12\times 10^{-10}$ Planck:2018vyg ; ParticleDataGroup:2024cfk , cannot be explained by the Standard Model (SM) of particle physics and hence provides evidence for new physics beyond the SM (BSM) Sakharov:1967dj . An attractive BSM scenario capable of dynamically generating the BAU in the early Universe consists of baryogenesis via leptogenesis Fukugita:1986hr , which relies on the interactions of right-handed neutrinos (RHNs) in the seesaw extension of the SM Minkowski:1977sc ; Yanagida:1979as ; Yanagida:1980xy ; Gell-Mann:1979vob ; Mohapatra:1979ia , thus linking the generation of the BAU to the origin of the tiny masses of the left-handed neutrinos in the SM Buchmuller:2005eh ; Chun:2017spz ; Bodeker:2020ghk . In its standard implementation, leptogenesis proceeds via the out-of-equilibrium decays of heavy RHNs, which simultaneously violate charge–parity invariance ($CP$) and baryon-minus-lepton number ($B\!-\!L$). These decays first produce a primordial lepton asymmetry, which is subsequently converted by means of the chemical transport in the thermal plasma, including electroweak sphalerons Kuzmin:1985mm , to a primordial baryon asymmetry.
The requirement of sufficient $CP$ violation in RHN decays bounds the RHN mass scale from below, $M_{N}\gtrsim 10^{9}\,\textrm{GeV}$. This constraint is known as the Davidson–Ibarra bound Davidson:2002qv ; Buchmuller:2002rq (see Ref. Garbrecht:2024xfs for a recent refinement). It assumes a hierarchical RHN mass spectrum (e.g., $M_{3}\gg M_{2}\gg M_{1}\equiv M_{N}$ in models with three RHNs) and can be circumvented in the case of a quasi-degenerate RHN mass spectrum (e.g., $M_{3}\gg M_{2}\simeq M_{1}$ or $M_{3}\simeq M_{2}\simeq M_{1}$), which gives rise to the scenario of resonant leptogenesis Pilaftsis:2003gt ; Pilaftsis:2005rv (see also Refs. Akhmedov:1998qx ; Klaric:2020phc ; Klaric:2021cpi ). In this paper, we will, however, ignore the possibility of resonantly enhanced $CP$ asymmetry in RHN decays and focus instead on an alternative route to low-scale leptogenesis at RHN masses $M_{N}\ll 10^{9}\,\textrm{GeV}$: wash-in leptogenesis Domcke:2020quw .
The key idea behind wash-in leptogenesis is to liberate the RHN decays from the requirement of sufficient $CP$ violation. Indeed, in wash-in leptogenesis, RHN interactions are treated as mere spectator processes and on the same footing as the electroweak sphalerons. From the perspective of wash-in leptogenesis, sphalerons and RHN decays represent nothing but sources of $B\!+\!L$ violation and $B\!-\!L$ violation, respectively. $CP$ violation, on the other hand, is attributed to new $CP$-violating dynamics at high temperatures, $T\gg M_{N}$, which set the required $CP$-violating initial conditions for wash-in leptogenesis. In this sense, wash-in leptogenesis does not represent a complete baryogenesis scenario on its own, but calls for an ultraviolet completion — a chargegenesis mechanism that is responsible for the generation of primordial $CP$-violating charge asymmetries at high energies. Wash-in leptogenesis then acts on the nontrivial configuration of chemical potentials induced by chargegenesis and reshuffles the primordial charge asymmetries in a $B\!-\!L$-violating manner. The SM notably offers more than ten conserved global charges at high temperatures that could be produced during chargegenesis Domcke:2020quw , which renders baryogenesis via leptogenesis via chargegenesis significantly more versatile than baryogenesis via leptogenesis. As a consequence, the range of possible chargegenesis scenarios is extremely broad; primordial charge asymmetries suitable for wash-in leptogenesis can, e.g., be produced during axion inflation Domcke:2022kfs , by evaporating primordial black holes Schmitz:2023pfy , in the decay of heavy Higgs doublets Mukaida:2024eqi , or via new nonminimal gravitational interactions Mojahed:2024yus .
As the temperature of the thermal bath drops, the last global charge to be erased by the SM interactions is the asymmetry between right-handed electrons and left-handed positrons, $q_{e}=n_{e}-\bar{n}_{e}$. Therefore, as long as the temperature scale of wash-in leptogenesis exceeds the equilibration temperature of the electron Yukawa interaction, $T_{e}\simeq 85\,\textrm{TeV}$ Bodeker:2019ajh , it is always possible to realize successful wash-in leptogenesis by assuming a nonzero primordial electron charge, $q_{e}^{\rm ini}\neq 0$. Correspondingly, the authors of Ref. Domcke:2020quw concluded that wash-in leptogenesis can take place at temperatures as low as $T\simeq 100\,\textrm{TeV}$. The purpose of the present paper is to make this statement more precise and study wash-in leptogenesis at $T\lesssim T_{e}$, where $q_{e}/T^{3}$ can no longer be treated as a constant, i.e., in a temperature regime where the asymmetry wash-in by RHN processes competes with the wash-out of the primordial $q_{e}$ input charge due to the progressing equilibration of the electron Yukawa interaction. To this end, we shall extend the treatment in Ref. Domcke:2020quw and supplement the Boltzmann equations for the three flavor asymmetries $q_{\Delta_{\alpha}}$, where $\Delta_{\alpha}=B/3-L_{\alpha}$ and $\alpha=e,\mu,\tau$, that are being produced during wash-in leptogenesis by a Boltzmann equation for the time-dependent electron charge $q_{e}$. Solving this coupled set of Boltzmann equations will then allow us to conclude that wash-in leptogenesis is able to produce the BAU for RHN masses as low as $M_{N}\simeq 7\,\textrm{TeV}$. In summary, in standard thermal leptogenesis where RHN decays act as a source of both $CP$ and $B\!-\!L$ violation, the RHN mass scale must satisfy the Davidson–Ibarra bound, $M_{N}\gtrsim 10^{9}\,\textrm{GeV}$; if RHN decays only act as a source of $B\!-\!L$ violation, it suffices to demand that $M_{N}\gtrsim 7\,\textrm{TeV}$.
Boltzmann equations — Let us begin by recalling a few key results from Ref. Domcke:2020quw . The evolution of the $q_{\Delta_{\alpha}}$ flavor asymmetries during $N_{1}$ leptogenesis is governed by Pilaftsis:2003gt ; Pilaftsis:2005rv
$-\left(\partial_{t}+3H\right)q_{\Delta_{\alpha}}=\varepsilon_{1\alpha}\,\Gamma_{1}\left(n_{N_{1}}-n_{N_{1}}^{\rm eq}\right)-\sum_{\beta}\gamma_{\alpha\beta}^{\rm w}\,\frac{\mu_{\ell_{\beta}}+\mu_{\phi}}{T}\,.$ (1)
Here, the negative sign on the left-hand side reflects the fact that the lepton flavor number $L_{\alpha}$ enters $\Delta_{\alpha}$ with a negative sign. The first term on the right-hand side is the source term for $q_{\Delta_{\alpha}}$ in standard thermal leptogenesis. In the following, we will assume vanishingly small $CP$ violation in the RHN sector, $\varepsilon_{1\alpha}\lll 1$, and simply ignore this term. Furthermore, the second term on the right-hand side is the wash-out term in standard thermal leptogenesis, where the total wash-out interaction density receives contributions from RHN inverse decays as well as from lepton-number violating ($\Delta L=2$) and lepton-flavor-violating ($\Delta L=0$) processes. This interaction density is multiplied by $\mu_{\ell_{\beta}}+\mu_{\phi}$, i.e., the sum of the chemical potentials of the SM charged-lepton and Higgs doublets. At $T\sim T_{e}$, the chemical potentials of the 16 SM particle multiplets, $\mu_{i}$ ($i=e,\mu,\tau,\ell_{e},\ell_{\mu},\ell_{\tau},u,c,t,d,s,b,Q_{1},Q_{2},Q_{3},\phi$), are subject to 16 constraints: 11 conditions for 11 linearly independent Yukawa and sphaleron interactions in chemical equilibrium and 5 conservation laws for 5 conserved global charges ($\Delta_{e,\mu,\tau}$, hypercharge $Y$, and the asymmetry in right-handed electrons). Solving this linear system of constraint equations for the 16 chemical potentials $\mu_{i}$, one finds Domcke:2020kcp ; Domcke:2020quw
$\begin{pmatrix}\mu_{\ell_{e}}+\mu_{\phi}\\ \mu_{\ell_{\mu}}+\mu_{\phi}\\ \mu_{\ell_{\tau}}+\mu_{\phi}\end{pmatrix}=\begin{pmatrix}-\frac{5}{13}\\ \frac{4}{37}\\ \frac{4}{37}\end{pmatrix}\,\bar{\mu}_{e}-\begin{pmatrix}\frac{6}{13}&0&0\\ 0&\frac{41}{111}&\frac{4}{111}\\ 0&\frac{4}{111}&\frac{41}{111}\end{pmatrix}\begin{pmatrix}\bar{\mu}_{\Delta_{e}}\\ \bar{\mu}_{\Delta_{\mu}}\\ \bar{\mu}_{\Delta_{\tau}}\end{pmatrix}\,,$ (2)
where the second term on the right-hand side contains the standard flavor coupling matrix Barbieri:1999ma ; Abada:2006fw ; Nardi:2006fx ; Abada:2006ea ; Blanchet:2006be ; Antusch:2006cw induced by the spectator processes in the thermal bath Buchmuller:2001sr ; Garbrecht:2014kda ; Garbrecht:2019zaa , while the first term is the relevant term for wash-in leptogenesis.
In presence of a primordial charge asymmetry, $\bar{\mu}_{e}\neq 0$, the processes contributing to the wash-out rate will not drive the $q_{\Delta_{\alpha}}$ to zero. Quite the opposite, these processes will now wash in nonzero lepton flavor asymmetries $q_{\Delta_{\alpha}}\equiv\nicefrac{{1}}{{6}}\>\bar{\mu}_{\Delta_{\alpha}}T^{2}$ proportional to $q_{e}\equiv\nicefrac{{1}}{{6}}\>\bar{\mu}_{e}T^{2}$, even if $q_{\Delta_{\alpha}}=0$ initially. Assuming $q_{e}/T^{3}=\textrm{const}$, wash-in leptogenesis in the strong wash-in regime (conventionally, strong wash-out regime) at $T\gtrsim T_{e}$ thus results in a $B\!-\!L$ asymmetry $q_{B-L}=-\nicefrac{{3}}{{10}}\,q_{e}$ Domcke:2020quw .
Around $T\sim T_{e}$, the assumption of constant $q_{e}/T^{3}$, however, begins to break down and $q_{e}/T^{3}$ needs to be described as a time-dependent quantity. The relevant Boltzmann equation for $q_{e}$ around the equilibration temperature of the electron Yukawa interaction has been derived in Ref. Bodeker:2019ajh ,
$\left(\partial_{t}+3H\right)q_{e}=-\Gamma_{e}\left[\frac{711}{481}\,q_{e}+\frac{5}{13}\,q_{\Delta_{e}}-\frac{4}{37}\left(q_{\Delta_{\mu}}+q_{\Delta_{\tau}}\right)\right]\,,$ (3)
where $\Gamma_{e}$ is the equilibration rate of right-handed electrons.
Let us now simplify the notation and bring the Boltzmann equations in Eqs. (1) and (3) into a more convenient form. First, we note that, around $T\sim T_{e}$, the standard wash-out rate is dominated by RHN inverse decays, such that
$\gamma_{\alpha\beta}^{\rm w}=p_{\alpha}\,\delta_{\alpha\beta}\,\frac{T^{3}}{6}\,\Gamma_{\rm w}\,,\qquad p_{\alpha}=\frac{\Gamma_{1\alpha}}{\Gamma_{1}}\,.$ (4)
$p_{\alpha}$ is the branching ratio for RHN decays into lepton flavor $\alpha$, and $\Gamma_{\rm w}$ is the rate of wash-out via inverse decays Buchmuller:2004nz ,
$\Gamma_{\rm w}=\frac{3g_{N}}{\pi^{2}}\,\frac{\widetilde{m}_{1}M^{2}}{8\pi v^{2}}\,z^{2}K_{1}\left(z\right)\,.$ (5)
Here, $g_{N}=2$ counts the RHN spin degrees of freedom (DOFs), $M$ denotes the RHN mass, $z=M/T$ functions as a time variable, $K_{1}$ is a modified Bessel function, $v=174\,\textrm{GeV}$ is the Higgs vacuum expectation value, and $\widetilde{m}_{1}$ is an effective light neutrino mass defined in terms of $M$ and the neutrino Dirac mass matrix $m_{D}$ in the seesaw model Plumacher:1996kc ,
$\widetilde{m}_{1}=\frac{\left(m_{D}^{\dagger}m_{D}^{\vphantom{\dagger}}\right)_{11}}{M_{1}}\,.$ (6)
Figure 1: Example solutions of the set of Boltzmann equations in Eqs. (13) to (16) in the weak / strong wash-in regime (top / bottom row). Solid / dashed lines denote positive / negative comoving asymmetries $Q$. The $Q_{B-L}$ values at $z\gg 1$ allow one to read off the efficiency factor defined in Eq. (23), $Q_{B-L}\left(z\gg 1\right)=C_{\rm win}\,\kappa$. The solutions in the first / second column correspond to positive / negative $\kappa$ values. In all four plots, the RHN branching ratios $p_{e,\mu,\tau}$ are set to the optimal values that maximize $\left|\kappa\right|$ for the given neutrino masses $\widetilde{m}_{1}$ and $M$.
Meanwhile, the electron equilibration rate can be written as $\Gamma_{e}=G_{e}\,T$, where the coefficient $G_{e}$ receives contributions from $2\rightarrow 2$ processes involving one electron Yukawa vertex as well as from $1\leftrightarrow 2$ decays and inverse decays of the SM Higgs bosons, right-handed electrons and left-handed leptons. In fact, the latter processes also receive corrections from $1n\leftrightarrow 2n$ scatterings with soft gauge boson exchanges, which can be taken care of by Landau–Pomeranchuk–Migdal (LPM) resummation Landau:1953um ; Landau:1953gr ; Migdal:1956tc . Taking all these processes into account, the authors of Ref. Bodeker:2019ajh find
$C_{e}=\frac{3h_{e}^{2}}{1024\pi}\Bigg{[}$ $h_{t}^{2}c_{t}+\left(3g^{2}+g^{\prime 2}\right)\left(c_{\ell}+\ln\frac{1}{3g^{2}+g^{\prime 2}}\right)$ (7)
$+4g^{\prime 2}\left(c_{e}+\ln\frac{1}{4g^{\prime 2}}\right)\Bigg{]}\,,$ (8)
where $c_{t}$, $c_{\ell}$, and $c_{e}$ receive contributions from both $2\rightarrow 2$ processes and LPM-resummed $1n\leftrightarrow 2n$ processes,
$c_{t}$ $=2.82+1.48\,,$ (9)
$c_{\ell}$ $=3.52+0.776\,,$ (10)
$c_{e}$ $=2.69+2.03\,.$ (11)
Furthermore, $h_{e}$ and $h_{t}$ in Eq. (7) denote the electron and top-quark Yukawa couplings, while $g$ and $g^{\prime}$ are the $SU(2)_{L}$ and $U(1)_{Y}$ gauge couplings. In our numerical analysis, we start from the SM values of all four couplings at the $Z$ pole and then use their one-loop renormalization group equations (RGEs) to evolve them up to the RGE scale $\mu=\pi T$.
Next, we introduce the comoving charge asymmetries in units of the initial electron asymmetry $q_{e}^{\rm ini}$ at $T_{\rm ini}\gg T_{e}$,
$Q_{\Delta_{\alpha}}=\left(\frac{a}{a_{\rm ini}}\right)^{3}\frac{q_{\Delta_{\alpha}}}{q_{e}^{\rm ini}}\,,\qquad Q_{e}=\left(\frac{a}{a_{\rm ini}}\right)^{3}\frac{q_{e}}{q_{e}^{\rm ini}}\,,$ (12)
where $a\propto z$ is the Friedmann–Lemaître–Robertson–Walker scale factor. Using our above results for the interaction rates $\Gamma_{\rm w}$ and $\Gamma_{e}$, we can then rewrite the Boltzmann equations in Eqs. (1) and (3) as a set of equations for $Q_{\Delta_{\alpha}}$ and $Q_{e}$,
$Q^{\prime}_{\Delta_{e}}$ $=\,p_{e}W\left[-\frac{5}{13}\,Q_{e}-\frac{6}{13}\,Q_{\Delta_{e}}\right]\,,$ (13)
$Q^{\prime}_{\Delta_{\mu}}$ $=p_{\mu}W\left[\frac{4}{37}\,Q_{e}-\frac{41}{111}\,Q_{\Delta_{\mu}}-\frac{4}{111}\,Q_{\Delta_{\tau}}\right]\,,$ (14)
$Q^{\prime}_{\Delta_{\tau}}$ $=\,p_{\tau}W\left[\frac{4}{37}\,Q_{e}-\frac{4}{111}\,Q_{\Delta_{\mu}}-\frac{41}{111}\,Q_{\Delta_{\tau}}\right]\,,$ (15)
$Q_{e}^{\prime}$ $=\quad-E\left[\frac{711}{481}\,Q_{e}+\frac{5}{13}\,Q_{\Delta_{e}}-\frac{4}{37}\left(Q_{\Delta_{\mu}}+Q_{\Delta_{\tau}}\right)\right]\,.$ (16)
Here, $W$ denotes the dimensionless wash-out rate,
$W=\frac{\Gamma_{\rm w}}{zH}=\frac{6}{\pi^{2}}\,\frac{\widetilde{m}_{1}}{m_{*}}\,z^{3}K_{1}\left(z\right)\,,$ (17)
while $E$ is the dimensionless electron equilibration rate,
$E=\frac{\Gamma_{e}}{zH}=C_{e}\,\frac{M_{*}}{M}\,.$ (18)
In both expressions, $H$ is the Hubble rate, which can be written as follows during the era of radiation domination,
$H=\frac{T^{2}}{M_{*}}=\frac{M^{2}}{z^{2}M_{*}}\,,$ (19)
where $M_{*}$ denotes the reduced Planck mass $M_{\rm Pl}$ weighted by a factor taking into account the effective number of relativistic DOFs in the SM at high temperatures, $g_{*}=\nicefrac{{427}}{{4}}$,
$M_{*}=\left(\frac{90}{\pi^{2}g_{*}}\right)^{1/2}M_{\rm Pl}\simeq 7.12\times 10^{17}\,\textrm{GeV}\,.$ (20)
The mass $m_{*}$ in Eq. (17), on the other hand, is defined as Buchmuller:2004nz
$m_{*}=\frac{8\pi v^{2}}{M_{*}}\simeq 1.07\,\textrm{meV}\,.$ (21)
These definitions conclude our derivation of the Boltzmann equations. As evident from Eqs. (13) to (16), wash-in leptogenesis at $T\sim T_{e}$ effectively depends on four parameters: (1) the effective neutrino mass $\widetilde{m}_{1}$, which controls the strength of asymmetry wash-in, (2) the RHN mass $M$, which determines the temperature scale of wash-in leptogenesis and hence the importance of $q_{e}$ wash-out, and (3 and 4) the branching ratios $p_{e}$ and $p_{\mu}$ ($p_{\tau}$ follows from $p_{e}+p_{\mu}+p_{\tau}=1$).
Numerical results — We are now ready to numerically solve the Boltzmann equations for $Q_{\Delta_{\alpha}}$ and $Q_{e}$. In doing so, we shall assume vanishing initial lepton flavor asymmetries and a nonvanishing primordial electron asymmetry (stemming from an unspecified chargegenesis mechanism),
$Q_{\Delta_{\alpha}}\left(z\ll 1\right)=0\,,\qquad Q_{e}\left(z\ll 1\right)=1\,.$ (22)
Four example solutions of our set of Boltzmann equations for these initial conditions are shown in Fig. 1. For a given numerical solution, we quantify the efficiency of wash-in leptogenesis, in relation to the optimal outcome mentioned above, $q_{B-L}=-\nicefrac{{3}}{{10}}\,q_{e}$, in terms of the efficiency factor
$\kappa=\frac{1}{\nicefrac{{3}}{{10}}}\left[Q_{\Delta_{e}}\left(z\gg 1\right)+Q_{\Delta_{\mu}}\left(z\gg 1\right)+Q_{\Delta_{\tau}}\left(z\gg 1\right)\right]\,.$ (23)
For given $\kappa=\kappa\left(\widetilde{m}_{1},M,p_{e},p_{\mu}\right)$ and primordial input charge $q_{e}^{\rm ini}\equiv\nicefrac{{1}}{{6}}\>\bar{\mu}_{e}^{\rm ini}T^{2}$, one thus obtains a final baryon asymmetry
$\eta_{B}^{0}\left(\widetilde{m}_{1},M,p_{e},p_{\mu},q_{e}^{\rm ini}\right)=C_{\rm sph}\,C_{\rm win}\,\frac{g_{*,s}^{0}}{g_{*,s}^{\rm SM}}\,\kappa\left.\frac{q_{e}^{\rm ini}}{n_{\gamma}}\right|_{\rm ini}\,,$ (24)
with sphaleron and wash-in leptogenesis conversion factors $C_{\rm sph}=\nicefrac{{12}}{{37}}$ Harvey:1990qw ; Laine:1999wv and $C_{\rm win}=\nicefrac{{3}}{{10}}$, SM entropy DOFs today and at high temperatures, $g_{*,s}^{0}=\nicefrac{{43}}{{11}}$ and $g_{*,s}^{\rm SM}=\nicefrac{{427}}{{4}}$, and photon number density $n_{\gamma}=\zeta(3)\,g_{\gamma}T^{3}/\pi^{2}$, where $g_{\gamma}=2$.
The allowed $q_{e}^{\rm ini}$ range is bounded from above by the requirement that the primordial chiral asymmetry stored in right-handed electrons must not trigger the chiral instability of the SM plasma Joyce:1997uy ; Boyarsky:2011uy ; Akamatsu:2013pjd ; Hirono:2015rla ; Yamamoto:2016xtu ; Rogachevskii:2017uyc ; Kamada:2018tcs . Indeed, $\bar{\mu}_{e}^{\rm ini}$ (i.e., $q_{e}^{\rm ini}$ rescaled so that it has units of a chemical potential) must satisfy Bodeker:2019ajh
$\left.\frac{\bar{\mu}_{e}}{T}\right|_{\rm ini}\lesssim 9.6\times 10^{-4}\,,$ (25)
which is equivalent to the condition $\left.\mu_{Y,5}/T\right|_{\rm ini}\lesssim 1.4\times 10^{-3}$, where $\mu_{Y,5}$ is the chiral chemical potential; before the onset of wash-in leptogenesis, we have $\mu_{Y,5}=\nicefrac{{711}}{{481}}\,\bar{\mu}_{e}$ Domcke:2022kfs ; Domcke:2022uue . If this upper bound is violated, the SM chiral plasma instability (CPI) will develop and convert the chiral charge stored in right-handed electrons into helical hypermagnetic fields. The helicity stored in these fields will then in turn lead to a drastic overproduction of baryon asymmetry Domcke:2022uue via the mechanism of baryogenesis via helicity decay Giovannini:1997eg ; Giovannini:1997gp ; Giovannini:1999wv ; Giovannini:1999by ; Bamba:2006km ; Bamba:2007hf ; Fujita:2016igl ; Kamada:2016eeb ; Kamada:2016cnb around the time of the electroweak phase transition.
The upper bound on the primordial input charge, together with the requirement of successful baryogenesis, $\eta_{B}^{0}=\eta_{B}^{\rm obs}$, implies a lower bound on the efficiency factor $\kappa$,
$\eta_{B}^{0}\simeq\eta_{B}^{\rm obs}\,\,\Bigg{(}\frac{\kappa}{2.6\times 10^{-4}}\Bigg{)}\left(\frac{\bar{\mu}_{e}/T}{9.6\times 10^{-4}}\right)_{\rm ini}\,.$ (26)
In the following, we are therefore interested in identifying the region of parameter space where $\kappa\gtrsim 2.6\times 10^{-4}$. To this end, we perform a numerical scan over the four-dimensional parameter space spanned by $\widetilde{m}_{1}$, $M$, $p_{e}$, and $p_{\mu}$. More precisely, at each point in the $\widetilde{m}_{1}$ – $M$ plane, we scan over $p_{e}\in[0,1]$ and $p_{\mu}\in[0,1-p_{e}]$ and identify the branching ratios that either maximize $\kappa$ or $-\kappa$. This analysis provides us with two extremal solutions for the efficiency factor, $\kappa_{+}>0$ and $\kappa_{-}<0$, in the $\widetilde{m}_{1}$ – $M$ plane; see Fig. 2. Both solutions are physically viable, as we can always choose the sign of the initial $q_{e}^{\rm ini}$ charge appropriately.
The contour plots of $\kappa_{\pm}$ in Fig. 2 represent the main result of our analysis. A first observation looking at Fig. 2 is that the $\kappa_{\pm}$ contours behave differently at small and large effective neutrino masses, which is reflected in the underlying optimal branching ratios. At $\widetilde{m}_{1}\ll 5\,\textrm{meV}$, maximal $\left|\kappa_{+}\right|$ ($\left|\kappa_{-}\right|$) is achieved for $p_{e}=0$ and $p_{\mu,\tau}=\nicefrac{{1}}{{2}}$ ($p_{e}=1$ and $p_{\mu,\tau}=0$). Typical solutions of the Boltzmann equations in this regime of parameter space are shown in the top row of Fig. 1. At $\widetilde{m}_{1}\gg 5\,\textrm{meV}$, on the other hand, $\left|\kappa_{+}\right|$ ($\left|\kappa_{-}\right|$) is maximized for $p_{e}\simeq 1-2.0\,\textrm{meV}/\widetilde{m}_{1}$ ($p_{e}\simeq 1.1\,\textrm{meV}/\widetilde{m}_{1}$) in combination with $p_{\mu,\tau}=(1-p_{e})/2$ ($p_{\mu}=0$ and $p_{\tau}=1-p_{e}$); see the bottom row of Fig. 1 for typical solutions of the Boltzmann equations.
Figure 2: $\kappa_{+}>0$ (solid red contours) and $\kappa_{-}<0$ (dashed blue contours) solutions for the efficiency factor, which assume optimal RHN branching ratios $p_{e,\mu,\tau}$, as functions of the neutrino masses $\widetilde{m}_{1}$ and $M$. The transparent bands in the zoomed-in portion on the right-hand side indicate the ranges $\left|\kappa_{\pm}\right|=2\cdots 4\times 10^{-4}$, with the central lines corresponding to the $\left|\kappa_{\pm}\right|=3\times 10^{-4}$ contours.
The $\left|\kappa_{\pm}\right|=3\times 10^{-4}$ contours, furthermore, allow us to estimate lower bounds on the RHN mass $M$. In the case of positive (negative) $\kappa$, optimally chosen branching ratios, an effective neutrino mass around $\widetilde{m}_{1}\sim 100\,\textrm{meV}$ ($\widetilde{m}_{1}\sim 1\,\textrm{meV}$), and maximally allowed primordial input charge, wash-in leptogenesis can still successfully operate down to RHN masses of around $M\sim 8\,\textrm{TeV}$ ($M\sim 7\,\textrm{TeV}$). In fact, since the upper limit in Eq. (25) is not a perfectly sharp threshold and we rather expect the CPI to become gradually more important around primordial input charges around the threshold value, we estimate the uncertainty of our RHN mass bound by allowing the lower bound on $\kappa$ to vary in the range from $\kappa=2\cdots 4\times 10^{-4}$. In this way, we arrive at lower bounds of
$M_{\rm min}^{+}=8.1_{-1.0}^{+0.8}\,\textrm{TeV}\,,\qquad M_{\rm min}^{-}=7.1_{-0.7}^{+0.9}\,\textrm{TeV}\,,$ (27)
where $M_{\rm min}^{\pm}$ refer to the $\kappa_{\pm}$ solutions in Fig. 2, respectively.
Conclusions — The low RHN mass bounds in Eq. (27) are a remarkable result, which highlights that, assuming a large primordial input charge near the CPI threshold, the viable RHN mass range for successful wash-in leptogenesis extends around one order of magnitude below the equilibration temperature of the electron Yukawa interaction, $M_{\rm min}\sim 0.1\,T_{e}$. Our new bound on the RHN mass scale is thus one order of magnitude lower than the naive value of $M_{\rm min}\sim 100\,\textrm{TeV}$ initially stated in Ref. Domcke:2020quw and roughly five orders of magnitude below the classic Davidson–Ibarra bound on the RHN mass in standard thermal leptogenesis. We thus arrive at our main conclusion, which we already anticipated in the introduction: if, in the context of baryogenesis via leptogenesis, RHN interactions only serve the purpose to violate $B\!-\!L$ and $CP$ violation is attributed to new chargegenesis dynamics at higher energies, large viable regions in the neutrino mass parameter space open up; see Fig. 2. The relaxation of the lower bound on the RHN mass scale $M_{N}$ in wash-in leptogenesis has notably important consequences for model building, as $M_{N}$ no longer needs to lie several orders of magnitude above the electroweak scale.
In closing, we remark that the RHN mass bounds in Eq. (27) are still subject to remaining theory uncertainties due to our rough treatment of the CPI. In our analysis, we assumed that no CPI will develop as long as the bound in Eq. (25) is satisfied, while baryon number overproduction via helicity decay at the electroweak scale is unavoidable if it is not satisfied. This treatment could be refined by explicitly including the effect of anomalous $q_{e}$ violation in the $q_{e}$ Boltzmann equation, in combination with a new transport equation for hypermagnetic helicity (e.g., along the lines of Appendix C in Ref. Domcke:2019mnd ). In general, a proper understanding of the onset of the CPI at large chiral charge $\mu_{Y,5}$, however, calls for dedicated analyses based on magnetohydrodynamical simulations of the chiral plasma at temperatures $T\sim T_{e}$. We leave such refinements for future work and conclude instead that wash-in leptogenesis provides an intriguing framework for the generation of the BAU that can successfully operate at RHN masses as low as $M_{N}\sim 7\,\textrm{TeV}$.
Acknowledgments — The authors would like to thank Dietrich Bödeker, Kohei Kamada and Kyohei Mukaida for helpful discussions and comments. M. A. M. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG) Collaborative Research Centre “Neutrinos and Dark Matter in Astro- and Particle Physics” (SFB 1258). The work of K. S. and D. W. was supported by DFG through Research Training Group (Graduiertenkolleg) 2149: Strong and Weak Interactions — from Hadrons to Dark Matter.